40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 296     Entry time: Mon Feb 4 22:01:57 2008
Author: John 
Type: Summary 
Category: LSC 
Subject: Fibres auxiliary locking - Fibers 
I managed to couple ~75% of the light transmitted from the y arm, through the fibre, back to the SP table. I hoped that this would be a good way to match the beam from the fibre into the arm. Still no flashes. It looks like the cameras just aren't sensitive enough.
ELOG V3.1.3-